Grade 11General Physics

a projectile is fired vertically from earth surface with an initial speed of 10 kilo metre per second neglecting atmospheric retardation how far above the surface of earth would the projectile go(radius of earth is equals to 6400 kilometre)

shlesha dubey
10 Years agoGrade 11

1 Answer

Vikas TU
9 Years ago
Maximum distance it can cover would be:
H = u^2/2g => (10*1000)^2/2g = 10^7/2 meter.
or
10^4/2 km.
or
5000 km.